OpenStudy (anonymous):

A mixture was found to contain 1.50g of NH4CL, 0.80 g of NaCl, and 1.20 g of SiO2. What is the percentage of sand in this mixture?

OpenStudy (anonymous):

can someone help me please?

OpenStudy (anonymous):

sum up all masses and then divide mass of SiO2 with that summed up mass

OpenStudy (anonymous):

11.5?

OpenStudy (anonymous):

1.50 +0.80= 2.3 2.3 / 1.20= 11.5

OpenStudy (anonymous):

1.50g of NH4CL, 0.80 g of NaCl, and 1.20 g of SiO2. 1,5 + 0,8 +1,20 = 3,5 1,2/3,5=0,3428=34,28%